Calculate the change in total internal energy for a system that releases 2.71 × 104 kJ of heat and does 5.80 × 104 kJ of work on the surroundings. Enter your answer in scientific notation.  

Δ U = .....×^10..... kJ
Answer #1

Since heat is released, it must be negative

Q = -2.71*10^4 KJ

Since word is done by system, it must be negative

W = -5.80*10^4 KJ

Now use:

Δ U = Q + W

= -2.71*10^4 KJ -5.80*10^4 KJ

= -8.51*10^4 KJ

Answer: -8.51*10^4 KJ